OpenStudy (anonymous):

Show all your work. 2a+3b=-1 3a+5b=-2

OpenStudy (anonymous):

here 2a+3b=-1 so 2a = -1-3b ( by transposing 3b to RHS) a = (-1-3b)/2 putting the value of a in the 2nd equaion : 3(-1-3b)/2 + 5b=-2 (-3-9b)/2 + 5b=-2 (-3-9b+10b)/2 = -2 (-3+b)/2=-2 (-3+b)=-4 b=-4+3 b=-1 here 2a-3=-1 2a=2 a=1